OpenStudy (anonymous):

You invest an initial $2,000 in an account that has an annual interest rate of 6%, compounded daily. How much money will you have in the account after 15 years? Round your answer to the nearest whole number

OpenStudy (anonymous):

What formula or formulae do you think might be appropriate here?

OpenStudy (anonymous):

t=10, so p(10)=430*1.009^10?

OpenStudy (anonymous):

Apply this formula: \[A= P\left(1 + \frac{r}{n}\right)^{nt}\] * A = final amount * P = principal amount (initial investment) * r = annual nominal interest rate (as a decimal, not in percentage) * n = number of times the interest is compounded per year * t = number of years

OpenStudy (anonymous):

do i get 1.06

OpenStudy (anonymous):

No, you may've mistyped something into your calculator.

OpenStudy (anonymous):

for the first step

OpenStudy (anonymous):

Oh, for \(1+\frac{r}{n}\)? Not quite. You have to do \(1+\frac{0.06}{365}\) since the interest is compounded every day and there are \(365\) days in the year.

hero (hero):

Looks like someone already helped you?